OpenStudy (anonymous):

how do you integrate xe^(2x)dx?

OpenStudy (anonymous):

integral from 0 to 2

OpenStudy (saifoo.khan):

Use LATE principle?

OpenStudy (anonymous):

oh im using integration by parts and i know after we find the simplest from we plus in the limits into x but im not sue why the answer is 1/4 (3e^4+1) where did they get the 1/4 th

OpenStudy (anonymous):

-1/4+x^2/2+2x^3/3+x^4/2+0(x^5)

zepdrix (zepdrix):

\[\Large \int\limits \color{royalblue}{x} e^{2x}\;dx\]So by parts huh? Ok so we want the blue x to disappear right? So we'll let that be our `u`. Do you understand how to integrate the e^(2x) to find your `v`?

zepdrix (zepdrix):

\[\Large \int\limits\limits \color{royalblue}{x} e^{2x}\;dx\quad=\quad \frac{1}{2}x e^{2x}-\frac{1}{2}\int\limits e^{2x}\;dx\]By parts gives us this, yes? :o

OpenStudy (anonymous):

yes ive gotten that so far

OpenStudy (anonymous):

plugging in the limits is the confusing part.

zepdrix (zepdrix):

Sorry was cooking din din D:

zepdrix (zepdrix):

So finishing up the integration gives us:\[\Large \frac{1}{2}x e^{2x}-\frac{1}{4}e^{2x}\]Factoring some stuff:\[\Large \frac{1}{4}e^{2x}\left(2x-1\right)\] And now we need to plug in some limits? Hmm

zepdrix (zepdrix):

So at the upper limit we get:\[\Large \frac{1}{4}e^{4}(4-1)\]And then at the lower limit:\[\Large \frac{1}{4}e^0(-1)\]

zepdrix (zepdrix):

Giving us a grand total offffff:\[\Large \frac{1}{4}e^{4}(4-1)\quad - \quad \left[\frac{1}{4}e^0(-1)\right]\]

zepdrix (zepdrix):

The step where I factored out all of the stuff: 1/4 and e^2x, was that confusing at all? Do you understand why we were left with 2x in the first term?

OpenStudy (anonymous):

the factoring was indeed confusing but if i stare at theses steps for a while im sure ill get the hang of it. :) thank you very much! its a grand help!

zepdrix (zepdrix):

Try to keep this little math tid bit in mind:\[\Large 2+\frac{1}{4}\]If we want to factor 1/4 out of each term, we're `dividing` each term by 1/4. Which means we're actually `multiplying` each term by 4 as we factor our 1/4 out, right? :) \[\Large \frac{1}{4}\left(8+1\right)\]

zepdrix (zepdrix):

That's just an example, not numbers from our actual problem. A little tricky ^^ just something to keep in mind.

OpenStudy (anonymous):

ah, i am starting to connect things. thank you again. its been a great help.